# MESTI-MEGANet: Micro-expression Spatio-Temporal Image and Micro-expression Gradient Attention Networks for Micro-expression Recognition

### v1

@@ II-B1 Gradient Attention Block
This block, illustrated in Figure 4, explicitly models horizontal and vertical intensity gradients to localize micro-expression regions. Given an input image $X\in R^{B\times C\times H\times W}$, horizontal and vertical gradients at the spatial location $(i,j)$ are computed as:
@@
This attention map emphasizes regions with significant intensity transitions critical for micro-expression analysis. Figure 5 illustrate the gradient attention map constructed from our proposed MESTI as input image and gradient attention block.
@@ III-A Experiment scenarios and objectives
Experiment 02: This experiment evaluates the performance of MEGANet in MER. MEGANet is used as a MER network using Dynamic Image input, and its performance is compared with other studies that utilized Dynamic Image. Additionally, the experiment extends to test MEGANet using MESTI as input, with three corresponding MER networks: VGG19, ResNet + Micro-attention, and MEGANet.
Experiment 03: This experiment assesses the overall effectiveness of the proposed MER method, combining MESTI and MEGANet. The results of this experiment are compared with previously published MER methods to demonstrate the superiority of the proposed approach.
@@ III-B1 Datasets
The experiments are conducted on two publicly available micro-expression recognition datasets, namely CASME II and SAMM, which are widely used as standard benchmarks for MER and for comparison with previous studies.
  • CASME II: Part of the Chinese Academy of Sciences Micro-expression (CASME) dataset series, which includes CASME [42], CASME II [35], and $\text{CASME}^{2}$ [43]. Among these, CASME II is the most widely used dataset for micro-expression recognition tasks due to its significant improvements in both quantity and quality compared to the original CASME. CASME II contains a higher number of samples with more refined annotations, making it a reliable benchmark for evaluating micro-expression recognition networks. In contrast, $\text{CASME}^{2}$ is specifically designed for micro-expression spotting rather than recognition. It contains samples with longer durations and a lower frame rate, making it more suitable for identifying the precise temporal locations of micro-expressions within video sequences. This distinction highlights CASME II’s relevance and dominance in the context of micro-expression recognition tasks. CASME II consists of 247 micro-expressions collected from 26 subjects. The dataset was recorded using a high-speed camera at 200 frames per second (fps), and the face regions were cropped to a resolution of 280 × 340 pixels. These high frame rates and precise annotations make CASME II particularly well-suited for capturing the subtle and fleeting nature of micro-expressions.
  • SAMM: Another widely used dataset in micro-expression recognition research. It comprises 159 micro-expression (ME) samples collected from 32 participants. The dataset was captured using a grayscale camera operating at 200 frames per second (fps) under controlled lighting conditions to prevent flickering. This high-speed recording ensures that subtle micro-expressions are accurately captured and preserved. A notable strength of SAMM lies in its ethnic diversity. Unlike earlier datasets, which often lacked representation, SAMM features participants from 13 different ethnic backgrounds. This diversity enhances the dataset’s robustness and generalizability, making it a valuable resource for developing and evaluating micro-expression recognition networks across varied populations.
Figure 8: Visualization of MESTI and corresponding Gradient Attention Map characterize each emotion of micro-expression (Best viewed in color)
TABLE I: Comparison of input modalities in Micro-Expression Recognition (MER) with Train-Test split protocol.
Input modality for MER Network
VGG19
(CASMEII - SAMM)
Resnet50
(CASMEII - SAMM)
EfficientnetB0
(CASMEII - SAMM)
Static Apex frame 50.00% - 43.75% 46.15% - 50.00% 34.62% - 43.75%
Dynamic Optical flow (Onset - Apex) 50.00% - 50.00% 46.15% - 37.50% 26.92% - 43.75%
Optical flow (Apex - Offset) 53.85% - 50.00% 34.62% - 43.75% 46.15% - 43.75%
Dynamic image 57.69% - 50.00% 53.85% - 37.50% 53.85% - 50.00%
Affective motion image 50.00% - 43.75% 53.85% - 50.00% 46.15% - 43.75%
Active image 48.00% - 57.14% 44.00% - 50.00% 52.00% - 48.00%
MESTI (ours) 73.08% - 62.5% 65.38% - 56.25% 61.54% - 50.00%
@@ III-B2 Data preprocessing
The original five emotion labels are retained as annotated in both datasets. Therefore, in the CASME II dataset, the micro-expression labels include disgust (60), happiness (33), other (102), repression (27), and surprise (25). The SAMM dataset consists of the following emotion labels: anger (57), happiness (26), contempt (12), surprise (15), and other (49).
Figure 9: Results of each subject on CASMEII dataset.
Figure 10: Results of each subject on SAMM dataset.
TABLE II: Comparison of state-of-the-art methods and input modalities for MER with LOSO protocol in CASMEII Dataset.
No Method Network architecture Input modality Category
Accuracy
1 MER-GCN [53] 3DCNN Action unit 5 42.03%
2 MER-GCN [53] GCN Action unit 5 42.71%
3 SelectiveCNN [45] CNN Difference image 5 47.3%
4 LEARNet [51] VGG19 Dynamic Image 5 51.02%
5 Capsnet [50] CapsNet Apex image 5 51.62%
6 Capsnet [50] CNN Apex image 5 55.28%
7 AffectiveNet [55] Affective net Affective-motion image 4 61.02%
8 Active Imaging [57] OrigiNet Active Image 5 62.09%
9 OFF-ApexNet [46] OFF-Apex Net Apex image 5 62.18%
10 LGCCon [54] VGG16 + GI path + LI path Apex image 5 65.02%
11 Micro-attention [48] Resnet + Micro-attention Apex image 5 65.9%
12 FeatRef [41] InceptionNet + FeatRef Optical Flow 5 68.38%
13 GEME [56] Dualstream-Multitask Dynamic Image 5 75.2%
14 AMAN [52] CNN + Attention Block Image sequence 5 75.4%
15 LEARNet [51] CNN-LEARNet Dynamic Image 5 76.57%
16 Ours MEGANet Dynamic Image 5 68.16%
17 Ours VGG19 MESTI 5 69.39%
18 Ours Resnet + Micro-attention [48] MESTI 5 71.02%
19 Ours MEGANet MESTI 5 78.78%
TABLE III: Comparison of state-of-the-art methods and input modalities for MER with LOSO protocol in SAMM Dataset.
No Method Network architecture Input modality Category
Accuracy
1 Active Imaging [57] OrigiNet Active Image 5 34.02%
2 LGCCon [54] VGG16 + GI path + LI path Apex image 5 40.9%
3 LEARNet [51] VGG19 Dynamic Image 5 43.23%
4 LEARNet [51] CNN-LEARNet Dynamic Image 5 44.85%
5 Micro-attention [48] Resnet + Micro-attention Apex image 5 48.5%
6 GEME [56] Dualstream-Multitask Dynamic Image 5 55.88%
7 AffectiveNet [55] Affective net Affective-motion image 4 58.12%
8 FeatRef [41] InceptionNet + FeatRef Optical Flow 5 60.13%
9 AMAN [52] CNN + Attention Block Image sequence 5 68.85%
10 Ours MEGANet Dynamic Image 5 63.97%
11 Ours VGG19 MESTI 5 60.29%
12 Ours Resnet + Micro-attention [48] MESTI 5 63.24%
13 Ours MEGANet MESTI 5 71.32%
@@ III-C2 Experiment 02
MEGANet is evaluated using two input modalities: Dynamic Image and MESTI, as both modalities aim to represent a video as a single image. For the Dynamic Image input, MEGANet is compared with typical architectures that use Dynamic Image and evaluated using the LOSO protocol to ensure fairness. Building on the previous experiment, MEGANet is then directly compared with VGG19 and ResNet + Micro-attention, also using the LOSO protocol for evaluation and the MESTI input.
@@ III-C4 Specific configuration and training methodology
  • Optimizer: The Adam optimizer was employed with a learning rate of and weight decay of to optimize the network.
@@
  • Metric: The primary evaluation metric is accuracy. In the context of the LOSO protocol, accuracy is calculated as Equation 15:
$Accuracy=\frac{\sum_{i=1}^{S}T_{i}}{\sum_{i=1}^{S}N_{i}}\times 100\%$ (15)
where $T_{i}$ and $N_{i}$ are the number of correctly predicted samples and the number of samples, respectively, when the samples of the $i$-th subject is used as the test set.
@@ III-D1 Visual representation
The visual results of MESTI and its corresponding Gradient Attention Map are shown in Figure 8 to observe how MESTI captures the characteristic features of each micro-expression emotion type and how the Gradient Attention Map highlights the regions of interest within MESTI. A key observation is that MESTI effectively captures and highlights the defining motion patterns of micro-expressions, making them perceptible to the human eye in a single image representation.
@@
More specifically, both MESTI and the Gradient Attention Map successfully depict the characteristic Action Units corresponding to different micro-expression emotions. For Disgust, the key motion regions primarily appear around the eyebrows, one side of the nose, and the corners of the mouth. Repression manifests as subtle downward movements on both sides of the mouth and the chin. Happiness is expressed by an upward motion at the corners of the mouth, whereas Surprise is predominantly reflected in eyebrow elevation and lower lip movement. These findings highlight the capability of MESTI to encode motion dynamics effectively in a compact and visually interpretable format.
III-D2 Experiment 01
@@
Table I summarizes the comparative performance of various input modalities in the micro-expression recognition (MER) task, evaluated using deep learning network on the CASMEII and SAMM datasets. The results consistently demonstrate that MESTI outperforms all other input modalities across the three widely used CNN architectures: VGG19, ResNet50, and EfficientNetB0. Specifically, MESTI achieves the highest accuracy of 73.08% on CASMEII and 62.5% on SAMM with VGG19, surpassing the second-best input modality (Dynamic Image) by 15.39% and 12.5%, respectively. This superior performance underscores MESTI’s effectiveness in capturing subtle motion features, which are crucial for micro-expression recognition.
The second-best performing input modalities vary across networks but generally include Dynamic Image, Optical Flow, and Apex Frame, which have been commonly used in previous MER research. For example, Dynamic Image achieves 57.69% (CASMEII) and 50.00% (SAMM) with VGG19, while Optical Flow (Apex-Offset) achieves 53.85% (CASMEII) and 50.00% (SAMM). However, these modalities fall short of MESTI’s performance due to their inability to fully capture the transient and subtle nature of micro-expressions. Static images like the Apex Frame, while computationally efficient, lack temporal information, leading to suboptimal accuracy (50.00% on CASMEII and 43.75% on SAMM with VGG19). Optical Flow, though effective in encoding motion, is susceptible to noise from lighting variations and head movements, particularly for low-amplitude micro-expressions.
@@
To further validate MESTI’s effectiveness, we investigated whether its superior performance was specific to our proposed pipeline or if it could enhance other established MER architectures. The original input modalities are replaced by two previously published works with MESTI: VGG19 (originally using Dynamic Image) and ResNet + Micro-Attention (originally using Apex Frame). For VGG19, replacing the input with MESTI improved recognition accuracy from 51.02% to 69.39% on CASMEII and from 43.23% to 60.29% on SAMM. Similarly, for ResNet + Micro-Attention, using MESTI as input improved accuracy from 65.90% to 71.02% on CASMEII and from 48.5% to 63.24% on SAMM. These results confirm that MESTI not only enhances our proposed network but also significantly improves the performance of other MER architectures, demonstrating its capability to effectively represent micro-expression dynamics in a single image.
III-D3 Experiment 02
Proposed MEGANet is evaluated by comparing it to other architectures when using MESTI as input. Figures 9 and 10 illustrate the LOSO validation results for individual subjects, confirming that MEGANet outperforms both VGG19 and ResNet + Micro-Attention when trained on MESTI. This indicates that the integration of MESTI with MEGANet’s Gradient Attention mechanism leads to a substantial performance gain. The results, shown in Row 19 of Table II and Row 13 of Table III, demonstrate that the combination of MEGANet + MESTI achieves state-of-the-art performance, reaching 78.78% accuracy on CASMEII and 71.32% on SAMM, surpassing previously published methods.
To further assess MEGANet’s adaptability, an additional experiment is conducted using Dynamic Image as input, comparing its performance with other Dynamic Image-based architectures. On the SAMM dataset, MEGANet achieved 63.97% accuracy, surpassing all other methods utilizing Dynamic Image as input. On CASMEII, MEGANet attained 68.16% accuracy, which, while slightly lower than LEARNet (76.57%) and GEME (75.2%), still demonstrated highly competitive performance. These results suggest that while MEGANet is optimized for MESTI, it remains highly effective even when using Dynamic Image. This underscores the robustness of MEGANet’s Gradient Attention mechanism, which enhances the extraction of micro-movements across different input representations.
III-D4 Experiment 03
The synergy between MESTI and MEGANet establishes a new benchmark in MER. As shown in Table II, the proposed method outperforms all prior works, including AMAN [52] (75.4%) and GEME [56] (75.2%), by 3.38% on CASMEII. On SAMM (Table III), it achieves 71.32%, surpassing the second-best method (AMAN [52]: 68.85%) by 2.47%. This success is attributed to two key factors:
@@ III-D4 Experiment 03
  • MEGANet’s Gradient Attention mechanism, which focuses on intensity transitions (Figure 5) while residual blocks model long-range dependencies (Figure 6).

### v2

@@ II-B1 Gradient Attention Block
This block, illustrated in Figure 4, explicitly models horizontal and vertical intensity gradients to localize ME regions. Given an input image $X\in R^{B\times C\times H\times W}$, horizontal and vertical gradients at the spatial location $(i,j)$ are computed as:
@@
This attention map emphasizes regions with significant intensity transitions critical for ME analysis. Figure 5 illustrate the gradient attention map constructed from our proposed MESTI as input image and gradient attention block.
@@ II-B2 Residual - Attention Block
Figure 8: Visualization of MESTI and corresponding Gradient Attention Map characterize each emotion of ME (Best viewed in color)
@@ III-A Experiment scenarios and objectives
Experiment 02: This experiment evaluates the performance of MEGANet in MER. And analysis the effective of key block proposed in MEGANet.
Experiment 03: This experiment assesses the overall effectiveness of the proposed MER method, combining MESTI and MEGANet. The results of this experiment are compared with recent SOTA methods to demonstrate the superiority of the proposed approach.
@@ III-B1 Datasets
The experiments are conducted on three publicly available ME recognition datasets, namely SMIC-HS[27], CASME II[45] and SAMM[50], which are widely used as standard benchmarks for MER and for comparison with previous studies.
@@ III-B2 Data preprocessing
To ensure fair comparison with prior studies, this work conducts both 3-class and 5-class evaluations. In the 3-class evaluation, three common categories across the datasets are considered: positive, negative, and surprise. For the 5-class evaluation, the original emotion annotations provided in the CASME II and SAMM datasets are retained. Specifically, the CASME II dataset comprises the ME labels disgust (60 samples), happiness (33), other (102), repression (27), and surprise (25). The SAMM dataset includes the labels anger (57 samples), happiness (26), contempt (12), surprise (15), and other (49).
@@ III-C1 Experiment 01
TABLE I: Comparison of input modalities in MER with Train–Test split protocol.
Input modality for MER VGG19 ResNet50 EfficientNet-B0
CASME II SAMM CASME II SAMM CASME II SAMM
Static
Apex frame 50.00% 43.75% 46.15% 50.00% 34.62% 43.75%
Dynamic
Optical flow (Onset–Apex) 50.00% 50.00% 46.15% 37.50% 26.92% 43.75%
Optical flow (Apex–Offset) 53.85% 50.00% 34.62% 43.75% 46.15% 43.75%
Dynamic image 57.69% 50.00% 53.85% 37.50% 53.85% 50.00%
Affective motion image 50.00% 43.75% 53.85% 50.00% 46.15% 43.75%
Active image 48.00% 57.14% 44.00% 50.00% 52.00% 48.00%
MESTI (ours) 73.08% 62.50% 65.38% 56.25% 61.54% 50.00%
TABLE II: MESTI Input representation in published MER networks with LOSO protocol compared with the original research.
Input Network CASME II SAMM
ACC ACC
Apex Image* Micro-attention [55] 65.9% 48.5%
MESTI Micro-attention [55] 71.02% 63.24%
Dynamic Image* VGG19 [63] 51.02% 43.23%
MESTI VGG19 [63] 69.39% 60.29%
    Results from the original research (baseline input); other rows use our MESTI representation.
@@ III-C2 Experiment 02
MEGANet is evaluated through an ablation study. Two ablation scenarios are conducted. In the first, individual components of MEGANet: the Gradient Attention Block and the Residual Attention Block are isolated and evaluated independently. In the second, the performance of MEGANet is assessed with respect to the varying number of Residual Attention Blocks to determine the most suitable configuration. Both ablations used the LOSO protocol for evaluation.
@@ III-C4 Specific configuration and training methodology
  • Optimizer: The Adam optimizer was employed with a learning rate of 1e31e-3 and weight decay of 1e41e-4 to optimize the network.
@@
  • Metric: The primary evaluation metric is accuracy, unweighted F1-score (UF1), unweighted average recall (UAR).
@@ III-D1 Visual representation
The visual results of MESTI and its corresponding Gradient Attention Map are shown in Figure 8 to observe how MESTI captures the characteristic features of each ME emotion type and how the Gradient Attention Map highlights the regions of interest within MESTI. A key observation is that MESTI effectively captures and highlights the defining motion patterns of MEs, making them perceptible to the human eye in a single image representation.
@@
More specifically, both MESTI and the Gradient Attention Map successfully depict the characteristic Action Units corresponding to different ME emotions. For Disgust, the key motion regions primarily appear around the eyebrows, one side of the nose, and the corners of the mouth. Repression manifests as subtle downward movements on both sides of the mouth and the chin. Happiness is expressed by an upward motion at the corners of the mouth, whereas Surprise is predominantly reflected in eyebrow elevation and lower lip movement. These findings highlight the capability of MESTI to encode motion dynamics effectively in a compact and visually interpretable format.
TABLE III: Comparison with recent SOTA methods in 3-class evaluation with LOSO protocol
Method Year CASME II SMIC-HS SAMM
UF1 UAR ACC UF1 UAR ACC UF1 UAR ACC
FeatRef [51] 2022 0.892 0.887 0.701 0.708 0.737 0.716
Dual-ATME [65] 2023 0.765 0.751 0.817 0.646 0.658 0.646 0.562 0.538 0.714
DS - 3DCNN [30] 2023 0.789 0.806 0.788 0.755 0.783 0.792
FRL-DGT [29] 2023 0.919 0.903 0.743 0.749 0.772 0.758
SelfME [26] 2023 0.908 0.929 0.697 0.701
Micron-BERT [25] 2023 0.903 0.891 0.855 0.838
MERASTC [61] 2023 0.933 0.950 0.790 0.862
GLEFFN [11] 2023 0.883 0.911 0.771 0.786
MCCA - VNet [42] 2024 0.915 0.923 0.816 0.811 0.883 0.871
ROI + WArcFace [44] 2025 0.924 0.910 0.940 0.811 0.819 0.818 0.787 0.785 0.862
SODA4MER [43] 2025 0.887 0.881 0.886 0.888
OFVIG-Net [59] 2025 0.713 0.720 0.644 0.640 0.607 0.579
MESTI-MEGANet 2025 0.913 0.929 0.932 0.917 0.924 92.68 0.890 0.914 0.918
  • Bold indicates the best result in each column; underline indicates the second-best result.
  • “–” denotes that the metric was not reported in the cited work.
TABLE IV: Comparison with recent SOTA methods in 5-class evaluation with LOSO protocol.
Method Year CASME II SAMM
UF1 UAR ACC UF1 UAR ACC
GEME [68] 2021 0.735 - 75.20 0.454 - 55.88
MER-Supcon [60] 2022 0.729 - 73.58 0.625 - 67.65
CMNet [56] 2023 0.740 - 78.05 0.772 - 78.68
C3DBed [49] 2023 0.752 - 77.64 0.722 - 75.73
KPCANet [35] 2023 0.659 - 70.46 0.522 - 63.83
JGULF [10] 2024 0.807 - 82.04 0.720 - 80.71
AU GCN [39] 2024 0.776 - 81.85 0.757 - 79.82
SODA4MER [43] 2025 0.814 - 84.18 0.789 - 80.30
LRT3O [38] 2025 0.791 - 81.78 0.757 - 80.15
MELLM [57] 2025 0.485 0.534 64.34 - - -
MESTI-MEGANet 2025 0.779 0.786 82.04 0.791 0.803 80.88
  • Bold indicates the best result in each column; underline indicates the second-best result.
  • “–” denotes that the metric was not reported in the cited work.
III-D2 MESTI Representation compared with other input modalities
@@
Table I summarizes the comparative performance of various input modalities in the ME recognition task, evaluated using deep learning network on the CASMEII and SAMM datasets. The results consistently demonstrate that MESTI outperforms all other input modalities across the three widely used CNN architectures: VGG19, ResNet50, and EfficientNetB0. Specifically, MESTI achieves the highest accuracy of 73.08% on CASMEII and 62.5% on SAMM with VGG19, surpassing the second-best input modality (Dynamic Image) by 15.39% and 12.5%, respectively. This superior performance underscores MESTI’s effectiveness in capturing subtle motion features, which are crucial for ME recognition.
@@
To further validate MESTI’s effectiveness, we investigated whether its superior performance was specific to our proposed pipeline or if it could enhance other established MER architectures. The original input modalities are replaced by two previously published works with MESTI: VGG19 (originally using Dynamic Image) and Micro-Attention (originally using Apex Frame). The results, presented in Table II, show that for VGG19, replacing the input with MESTI improved recognition accuracy from 51.02% to 69.39% on CASMEII and from 43.23% to 60.29% on SAMM. Similarly, for Micro-Attention, using MESTI as input improved accuracy from 65.90% to 71.02% on CASMEII and from 48.5% to 63.24% on SAMM. These results confirm that MESTI not only enhances our proposed network but also significantly improves the performance of other MER architectures, demonstrating its capability to effectively represent ME dynamics in a single image.
TABLE V: Ablation study of key blocks of MEGANet.
Gradient Attention Block Residual Block Self-attention UF1 UAR ACC
0.788 0.821 80.49
0.746 0.775 75.61
0.8304 0.8609 83.54
0.8084 0.8551 81.10
0.917 0.924 92.68
TABLE VI: Ablation study of number of Residual Attention Blocks
Residual Attention Block UF1 UAR ACC
$\times$ 2 0.802 0.842 82.32
$\times$ 3 0.917 0.924 92.68
$\times$ 4 0.844 0.878 85.98
III-D3 Compared with State-of-the-art methods in MER
The comparative results with recent state-of-the-art methods are reported in Table III (for the 3-class evaluation) and Table IV (for the 5-class evaluation). Overall, the proposed method outperforms existing SOTA approaches on the SAMM and SMIC-HS datasets and achieves competitive performance on the CASME II dataset, as reflected across all three evaluation metrics: accuracy, UF1, and UAR.
3-class evaluation
On SMIC-HS, MESTI-MEGANet attains the best performance on all three metrics (UF1=0.917, UAR=0.924, ACC=92.68%). The accuracy margin over the next best method (ROI+WArcFace, ACC=0.818) is $\approx$ +10.9 percentage points, indicating a substantial gain. On SAMM, our method again ranks first across metrics (UF1=0.890, UAR=0.914, ACC=0.918), with an accuracy improvement of +5.6% over the strongest competitor (ROI+WArcFace, ACC=0.862). On CASMEII, our scores are competitive but not leading: UF1=0.913 and UAR=0.929, ACC=0.932. The best UF1/UAR are achieved by MERASTC (UF1=0.933, UAR=0.950), while the highest accuracy belongs to ROI+WArcFace (ACC=0.940). The gaps to the leaders are modest: 0.020 in UF1, 0.021 in UAR, and 0.008 in ACC.
5-class evaluation
On CASMEII, SODA4MER yields the best UF1 and ACC (UF1=0.814, ACC=84.18%). MESTI-MEGANet reaches UF1=0.779, UAR=0.786, ACC=82.04 and matches JGULF on accuracy (82.04), trailing SODA4MER by 2.14. On SAMM, MESTI-MEGANet attains the top results across all reported metrics (UF1=0.791, UAR=0.803, ACC=80.88). The gains are small but consistent: UF1 is slightly higher than SODA4MER (0.789), and accuracy exceeds JGULF (80.71) by 0.17 and SODA4MER (80.30) by 0.58. Note that UAR on SAMM is not commonly reported by most baselines, so direct UAR comparisons are limited.
Across protocols and datasets, MESTI-MEGANet delivers SOTA on SMIC-HS (3-class) and strong SOTA on SAMM (both 3-class and 5-class), while remaining competitive on CASMEII (second-best accuracy in 3-class; tied for accuracy but below the best UF1 in 5-class). These outcomes indicate that the method generalizes well to different datasets and label granularities, with the largest margins observed on SMIC-HS (dataset without apex frame annotated).
This success is attributed to two key factors:
@@ III-D3 Compared with State-of-the-art methods in MER
  • MEGANet’s Gradient Attention mechanism, which focuses on intensity transitions (Figure 5) while Residual Attention blocks model long-range dependencies (Figure 6).
III-D4 Approach with apex frame free dataset
In the SMIC dataset, apex frame annotations are not provided; hence, the apex frame information cannot be directly utilized to construct MESTI. As an alternative, in this study we adopt a simple strategy of selecting the middle frame.
Interestingly, the results on SMIC demonstrate strong performance despite the absence of apex frame supervision. This finding suggests that the proposed MESTI approach can remain effective even without precise apex frame information, highlighting its robustness and applicability in more challenging scenarios where apex annotations are unavailable.
III-E Ablation study
To evaluate the contribution of each component in MEGANet, we conducted an ablation study focusing on both the key building blocks and the number of Residual Attention Blocks.
As shown in Table V, removing any of the major components leads to a clear performance drop. Using only the Residual Block without Gradient Attention or Self-attention yields the lowest performance (UF1 = 0.746, UAR = 0.775, ACC = 75.61). Incorporating Self-attention alone provides some improvement (ACC = 83.54), while Gradient Attention Block alone achieves ACC = 81.10. The best performance is obtained when all three modules are integrated, resulting in significant gains (UF1 = 0.917, UAR = 0.924, ACC = 92.68). This demonstrates that the Gradient Attention Block, Residual Block, and Self-attention contribute complementary benefits, and their combination is essential for maximizing recognition accuracy.
In addition, Table VI investigates the impact of varying the number of Residual Attention Blocks. With two blocks, the model achieves ACC = 82.32, which increases substantially to 92.68 when three blocks are employed. Interestingly, adding a fourth block slightly reduces performance (ACC = 85.98), indicating potential overfitting or redundancy. These findings suggest that three Residual Attention Blocks provide the optimal balance between model complexity and representational power.